# Improving Medical Visual Representation Learning with Pathological-level Cross-Modal Alignment and Correlation Exploration

### v1

@@ Abstract
Learning medical visual representations from image-report pairs through joint learning has garnered increasing research attention due to its potential to alleviate the data scarcity problem in the medical domain. The primary challenges stem from the lengthy reports that feature complex discourse relations and semantic pathologies. Previous works have predominantly focused on instance-wise or token-wise cross-modal alignment, often neglecting the importance of pathological-level consistency. This paper presents a novel framework PLACE that promotes the Pathological-Level Alignment and enriches the fine-grained details via Correlation Exploration without additional human annotations. Specifically, we propose a novel pathological-level cross-modal alignment (PCMA) approach to maximize the consistency of pathology observations from both images and reports. To facilitate this, a Visual Pathology Observation Extractor is introduced to extract visual pathological observation representations from localized tokens. The PCMA module operates independently of any external disease annotations, enhancing the generalizability and robustness of our methods. Furthermore, we design a proxy task that enforces the model to identify correlations among image patches, thereby enriching the fine-grained details crucial for various downstream tasks. Experimental results demonstrate that our proposed framework achieves new state-of-the-art performance on multiple downstream tasks, including classification, image-to-text retrieval, semantic segmentation, object detection and report generation.
@@ 1 Introduction
Figure 1: The illustration of (a) the token-wise alignment, (b) clustered-based pseudo-pathology level and (c) our proposed pathological-level alignment. V-PORs are the abbreviation of visual pathology observation representations associated with specific anatomical regions.
@@
Some approaches have been proposed to promote the medical vision-language pre-training. For example, several studies [3, 6, 2] adopt the contrastive-learning-based local alignment between image patches and words as illustrated in Figure 1 (a). Nonetheless, applying the local alignment on the non-discriminative image patches and unimportant words introduces noisy information. Also, these token-wise alignment methods struggle to consider the pathological-level semantic information since this per-token-level granularity, e.g., one patch or word, is too small to represent pathology-level information. A group of studies [7, 6] improve the disease-level consistency by leveraging valuable disease labels, which however, are difficult to obtain in the real-world applications. To promote disease-aware alignment, [8] proposes a disease-level cross-modal alignment [8][9] further extend [8] by further distilling knowledge information into the disease-level cross-modal alignment [9]. Although improvements have been demonstrated, these so-called disease-level alignments are conducted on clustered representations as depicted in Figure 1 (b). Hence, they are less effectively applied to pathology alignment and might not adhere closely to a pathological-level alignment.
@@
    We demonstrate our framework to yield new state-of-the-art performance on a variety of downstream tasks: medical classification, object detection, semantic segmentation and zero-shot classification and image-to-text retrieval.
@@ 2.1 Medical Image and Report Pre-training
In recent years, great effort [10, 11, 12, 13, 14, 15, 16, 17] has been made to explore image-text joint training. By enforcing a cross-modal interaction and alignment, the learnt representation can demonstrate a potent generalisation capability that can be transferred to various multimodal tasks. This image-text joint training also serves to enhance visual representations with the incorporation of text, thereby playing a pivotal role in the medical domain where having a large-scale labelled dataset to train a single model is not only labour-intensive but also requires specialized expertise.
@@ 3 Method
This section presents the details of our proposed methods. Typically, the vision-language joint learning task endeavours to acquire a joint distribution $P(X,R)$ between the medical image set $X=\{x_{1},x_{2}...,x_{N}\}$ and the corresponding report set $R=\{r_{1},r_{2},...,r_{N}\}$. Each sample $e_{i}=<x_{i},r_{i}>$ in our context constitutes a medical image-report pair. With minimal training or a zero-shot scenario, a well-developed medical visual representation can be effectively adapted to various downstream tasks and exhibits favourable performance.
@@ 3.2 Image and Report Representation
Given an image-report pair $<x_{i},r_{i}>$, the first step is to obtain the representations for both the image and the report. Specifically, an image encoder $f_{I}$, e.g., ResNet50 [29] or ViT [30], transforms the image $x_{i}$ into the subregion representation, which is then flattened to a sequence of local visual tokens formed as $O^{LI}_{i,j}$. $i$ and $j$ denote $i^{th}$ sample in the batch and the $j^{th}$ patch in the image. Thereafter, average pooling is applied over the local visual tokens to derive the global image representation $O^{GI}_{i}$. $L$ and $G$ designate the global or local representation. Similarly, the report $r_{i}$ is embedded and encoded into a sequence of local textual tokens $U^{LR}_{i,j}$ by a text encoder $f_{R}$. $j$ refers to the $j^{th}$ token in the report. We follow the common practice of considering the $[CLS]$ token as the global report representation denoted by $U^{GR}_{i}$.
@@ 3.3 Instance-level Cross-modal Alignment
Images and their corresponding reports typically exhibit a substantial semantic congruity. The instance-level cross-modal alignment (ICMA) seeks to draw paired samples together while distancing the unpaired samples within the latent space at a global image-report level. This is accomplished by maximizing the mutual information between the image-report pairs, as articulated by Equation 1.
@@ 3.4 Pathological-level Cross-modal Alignment
An objective of cross-modal alignment at the instance level has demonstrated effectiveness in the acquisition of joint representations within the domain of natural scene. However, as a characteristic of medical VLP, distinct image-report pairs can demonstrate significant semantic similarity due to the subtle differences among images and the high similarities among reports. Models trained solely on instance-level contrastive loss struggle to learn the meaningful representation from samples sharing high similarities such as similar images and reports. Furthermore, various downstream medical tasks are increasingly reliant on a clinically accurate representation.
To this end, we propose a novel Pathological-level Cross-modal Alignment (PCMA) module that promotes consistency among image-report pairs in pathological observations. Our proposed PCMA module operates at a finer subject-level within each sample. Specifically, the PCMA module is developed to bring the anatomical region representation closer to its associated textual representation while creating distance between unpaired anatomical regions and sentences. This contrastive learning process is carried out with each sample. Consequently, our proposed PCMA module is less likely to be affected by high similarities among different samples, and encourages the model to delve deeper into the pathological structure within the samples.
@@
After having the textual pathological observation representation on hand, the next step is to acquire the visual pathological observation representation (V-POR). However, unlike the report, deriving the V-POR without accurate bounding box annotations and disease labels would be of great difficulty. Motivated by [10] which demonstrates that images of any resolution can be transformed to a fixed set of visual tokens while maintaining the semantic information, we propose to extract V-POR by exploiting the intrinsic structure of the image-report pair and by the aid of the T-POR without extra human annotations. Specifically, we first introduce a visual pathology observation extractor (VPOE) which is a transformer-based architecture (adhering to a Self-Attention $\rightarrow$ Cross-Attention structure) with a sequence of learnable pathology query tokens $Q\in\mathbb{R}^{N_{q}\times D}$. $N_{q}$ denotes the number of query tokens. Query tokens, designed to extract specific V-PORs from localized visual tokens, first performs a self-attention mechanism to capture contextual relationships within themselves. Then, they interact with these localized visual tokens through a cross-attention mechanism within the VPOE to align/extract the VPOs. Here, the query tokens serve as queries, while the localized visual tokens act as keys and values. The proposed PCMA module utilizes T-POR and contrastive learning to guide the model in exploiting the inherent structure of the image-report pair. This supervision aids in training the pathology query tokens to extract the most pertinent V-POR. Given the local visual tokens $U^{LI}_{i}$ and query tokens $Q$, the process to obtain the V-POR for $l^{th}$ transformer layer $O^{PI}_{i,l}$ in VDE is summarized by Equation 5. The upper right part of Figure 2 also illustrate this process.
@@
where $\mathrm{-Attn}$ refers to a vanilla attention mechanism and $Q_{0}=Q$. We consider the output of the last transformer layer in $VDE$ as the final visual pathological observation representation denoted by $O^{PI}_{i}$. After that, we enforce the alignment between the visual and textual POR. Nevertheless, there are no ground truth matching annotations between the visual and textual POR since the T-POR for one anatomical region can occur at any position in the report. Similarly to [2, 3], we adopt a cross-attention mechanism to investigate the matching between visual and textual POR. Specifically, for T-POR $O^{PI}_{i,j}$, we generate its corresponding cross-modal attended representation $C^{PI}_{i,j}$ via:
@@
Then, similar to ICMA, we apply the symmetric InfoNCE loss to maximise mutual information between VTOR $O^{PI}_{i}$ and its corresponding visual cross-modal attended representations (V-CARs) $C^{PI}_{i}$. Pathological-level report-to-image alignment loss $\mathcal{L}^{I\leftarrow R}_{PCMA}$ is formulated as:
@@
$w^{\prime}_{i}=\frac{w_{i}}{\sum_{k}{w_{k}}},\quad\quad w_{i}=\sum_{j\in sent_{i}}{a_{j}}$ (9)
@@ 3.5 Cross-modal Correlation Exploration
where $\bar{e}^{i}$ and $\bar{e}^{j}$ are the mean of $e^{i}$ and $e^{j}$ respectively. By calculating the covariance among each pair of patches (variables), we obtain a covariance matrix $\bm{\Sigma}^{i}\in\mathbb{R}^{N_{p}\times N_{p}}$ describing the correlation among different image patches in the $i^{th}$ sample. After that, to enforce the model to capture further fine-grain details and promote the cross-modal representation learning, we design a novel proxy task which requires it to predict this covariance matrix conditioned on the global report representation. We simply employ one linear layer as the CCP head to predict the covariance matrix: $\bm{H}^{i}=W_{h}\cdot\bm{U}^{GR}_{i}$ where $\bm{H}^{i}$ refers to the predicted covariance matrix for $i^{th}$ sample. After that, the Mean Squared Error (MSE) loss is used to supervise the learning of the cross-modal correlation prediction:
@@
The proposed objective CCE necessitates a significantly more fine-grained details understanding compared to the traditional MIM since it demands not only that the model comprehends the content within each image region from reports but also discerns the potential correlations among various image regions.
@@
Table 1: Results of semantic segmentation and image classification in the setting of zero-shot (for classification), 1%, 10% and 100% training samples. The evaluation metric for CheXpert is AUC. The best results are highlighted in red and the suboptimal methods are marked in blue respectively.
@@ 3.6 Overall Objective
Our proposed model is jointly trained with the three modules, i.e. the ICMA, PCMA and CCP modules, enforcing the framework to learn a more fine-grained, generalizable and pathologically discriminative medical visual representation. The overall training objectives is formulated as:
@@ 4.1 Pre-training Setup
Dataset: We follow most previous studies [2, 3, 8] to pre-train our proposed PLACE on the MIMIC-CXR dataset [36] and adopt the same data pre-processing procedure as [8, 9]. The images of the lateral view are removed since the downstream tasks only contain the frontal view images. Reports are formed by concatenating the Finding and Impression sections. We remove samples with an empty report or less than three words, resulting in a roughly total of 217,000 image-report pairs.
Implementation Details: Following [2, 3, 8], we employ ResNet-50 [29] and ClinicalBERT [37] as the backbone of our image and report encoder. Note that our proposed method is model-agnostic and can be applied to various backbones such as vision transformer and convolutional networks. We also report the results of adopting the vision transformer as the image encoder in classification tasks. Images are first resized to $256\times 256$ while maintaining the original size ratio with zero-padding for the smaller dimension, and then randomly cropped to $224\times 224$ during training. We adopt the AdamW [38] as the optimizer with a learning rate of $44e-$ 4 and weight decay of $25e-$ 2. The batch size is set to 128 on three A100-40G GPU cards. We train our model for $50$ epochs with an early stop mechanism that terminates the training without seeing a decrease in validation loss for more than $10$ epochs. Consistent with [8, 39], we set the softmax temperature $\tau_{1}$ and $\tau_{2}$ to $0.07$ and $0.10$ respectively. The split patch size $PS$ in CCP module is set to $32$ (determined by a small grid search in $\{16,28,32,56\}$, resulting in a $N_{p}$ of $49$. We set the number of pathological query tokens in Equation 8 to $12$. The loss weights $\lambda$ and $\beta$ are set to $0.5$ determined by a small grid search in $\{0.1,0.25,0.5,1\}$.
@@ 4.2 Downstream Tasks
Medical Semantic Segmentation. The SIIM Pneumothorax [40] and RSNA Pneumonia [41] datasets are adopted to assess the capability of our model for this task. Consistent with most previous works, we adopt the U-Net architecture and employ our pre-trained image encoder as the the encoder backbone (weight frozen), while fine-tuning the decoder using 1%, 10% and 100% training samples. Dice score [42] is selected as the evaluation metric.
@@
Medical Image Classification. We verify the effectiveness of our pre-trained image encoder for medical image classification on COVIDx [44] and CheXpert [45] datasets. Following prior work [8, 9, 3], we adopt linear probing which freezes the pre-trained image encoder and only trains the classification head on COVIDx dataset on three scenarios, i.e., 1%, 10% and 100% training samples. Additionally, we also gauge the zero-shot generalization capability of our model on CheXpert and COVIDx datasets. Same as  [8, 9], we also report the results of taking the vanilla ViT [30] as image encoder for classification tasks.
@@
Medical Report Generation. We further investigate PLACE’s ability to comprehend cross-modal information on the MIMIC-CXR dataset through the task of report generation. We adopt two fundamental and typical architecture as our baseline (1) ST [46]: a visual extractor with transformer-based encoder-decoder architecture and (2) C2GPT2 [47]: a vision encoder-language decoder architecture. Note that the baselines are standard transformer-based models without any advanced techniques or additional data sources tailored specifically to MRG. In our approach, we employ the pretrained image encoder from PLACE as the visual extractor in ST and as the vision encoder in Res2GPT2, maintaining fixed weights while only optimizing other model components. The official data split is adopted. We evaluable the performance of the model by the commonly utilized metrics including BLEU [48], METEOR [49], ROUGE-L [50] and CIDEr [51].
@@
Prec @@@@ 1 Prec @@@@ 2 Prec @@@@ 5 Prec @@@@ 10

### v2

@@ Abstract
Learning medical visual representations from image-report pairs through joint learning has garnered increasing research attention due to its potential for transferring acquired knowledge to various downstream medical tasks. Previous works have predominantly focused on instance-wise or token-wise cross-modal alignment, often neglecting the importance of pathological-level consistency. This paper presents a novel framework PLACE that promotes the Pathological-Level Alignment and enriches the fine-grained details via Correlation Exploration without additional human annotations. Specifically, we propose a novel pathological-level cross-modal alignment (PCMA) approach to maximize the consistency of pathology observations from both images and reports. To facilitate this, a Visual Pathology Observation Extractor is introduced to extract visual pathological observation representations from localized tokens. The PCMA module operates independently of any external disease annotations, enhancing the generalizability and robustness of our methods. Furthermore, we design a proxy task that enforces the model to identify correlations among image patches, thereby enriching the fine-grained details crucial for various downstream tasks. Experimental results demonstrate that our proposed framework achieves new state-of-the-art performance on multiple downstream tasks, including classification, image-to-text retrieval, semantic segmentation, object detection and report generation. Code is available at https://github.com/Markin-Wang/PLACE.
@@ 1 Introduction
Figure 1: The illustration of (a) the token-wise alignment, (b) clustered-based pseudo-pathology level and (c) our proposed pathological-level alignment. V-PORs is the abbreviation of visual pathology observation representations which are associated with specific anatomical regions.
@@
Some approaches have been proposed to promote the medical vision-language pre-training. For example, several studies [3, 6, 2] adopt the contrastive-learning-based local alignment between image patches and words as illustrated in Figure 1 (a). Nonetheless, applying the local alignment on non-discriminative image patches and unimportant words introduces noisy information. Also, these token-wise alignment methods struggle to consider the pathological-level semantic information since this per-token-level granularity, e.g., one patch or word, is too small to represent pathology-level information. A group of studies [7, 6] improve the disease-level consistency by leveraging valuable disease labels, which however, are difficult to obtain in the real-world applications. To promote disease-aware alignment, [8] proposes a disease-level cross-modal alignment [8][9] further extend [8] by further distilling knowledge information into the disease-level cross-modal alignment. Although improvements have been demonstrated, these so-called disease-level alignments are conducted on clustered representations as depicted in Figure 1 (b). Hence, they are less effectively applied to pathology alignment and might not adhere closely to a pathological-level alignment.
@@
    We demonstrate our framework to yield new state-of-the-art performance on a variety of downstream tasks: medical classification, object detection, semantic segmentation, and image-to-text retrieval.
@@ 2.1 Medical Image and Report Pre-training
In recent years, great effort [10, 11, 12, 13, 14, 15, 16, 17] has been made to explore image-text joint training. By enforcing a cross-modal alignment, the learnt representation shows a potent generalisation capability that can be transferred to various multimodal tasks. This image-text joint training also serves to enhance visual representations with the incorporation of text, thereby playing a pivotal role in the medical domain where having a large-scale labelled dataset to train a single model is not only labour-intensive but also requires specialized expertise.
@@ 3 Method
This section presents the details of our proposed methods. Typically, the vision-language joint learning task endeavours to acquire a joint distribution $P(X,R)$ between the medical image set $X=\{x_{1},x_{2}...,x_{N}\}$ and the corresponding report set $R=\{r_{1},r_{2},...,r_{N}\}$. Each sample $e_{i}=<x_{i},r_{i}>$ in our context constitutes a medical image-report pair. With minimal training or a zero-shot scenario, a well-developed medical visual representation can be effectively adapted to various downstream tasks and exhibit favourable performance.
@@ 3.2 Image and Report Representation
Given an image-report pair $x_{i},r_{i}>$, the first step is to obtain the representations for both the image and the report. Specifically, an image encoder $f_{I}$, e.g., ResNet50 [29] or ViT [30], transforms the image $x_{i}$ into a subregion representation, which is then flattened to a sequence of local visual tokens formed as $O^{LI}_{i,j}$. $i$ and $j$ denote $i^{th}$ sample in the batch and the $j^{th}$ patch in the image. Thereafter, average pooling is applied over the local visual tokens to derive the global image representation $O^{GI}_{i}$. $L$ and $G$ designate the global or local representation. Similarly, the report $r_{i}$ is embedded and encoded into a sequence of local textual tokens $U^{LR}_{i,j}$ by a text encoder $f_{R}$. $j$ refers to the $j^{th}$ token in the report. We follow the common practice of considering the $[CLS]$ token as the global report representation denoted by $U^{GR}_{i}$.
@@ 3.3 Instance-level Cross-modal Alignment
Images and their corresponding reports typically exhibit a substantial semantic congruity. The instance-level cross-modal alignment (ICMA) stage seeks to draw paired samples together while distancing the unpaired samples within the latent space at a global image-report level. This is accomplished by maximizing the mutual information between the image-report pairs, as articulated by Equation 1.
@@ 3.4 Pathological-level Cross-modal Alignment
An objective of cross-modal alignment at the instance level has demonstrated effectiveness in the acquisition of joint representations within the domain of natural scenes. However, as a characteristic of medical VLP, distinct image-report pairs can demonstrate significant semantic similarity due to the subtle differences among images and the high similarities among reports. Models trained solely on instance-level contrastive loss struggle to learn a meaningful representation from samples sharing high similarities such as similar images and reports. Furthermore, various downstream medical tasks are increasingly reliant on a clinically accurate representation.
To this end, we propose a novel Pathological-level Cross-modal Alignment (PCMA) module that promotes consistency among image-report pairs in pathological observations. The PCMA module operates at a finer subject-level within each sample. Specifically, it is developed to bring the anatomical region representation closer to its associated textual representation while creating distance between unpaired anatomical regions and sentences. This contrastive learning process is carried out with each sample. Consequently, the PCMA module is less likely to be affected by high similarities among different samples, and encourages the model to delve deeper into the pathological structure within the samples.
@@
After having the textual pathological observation representation on hand, the next step is to acquire the visual pathological observation representation (V-POR). However, unlike the report, deriving the V-POR without accurate bounding box annotations and disease labels would be of great difficulty. Motivated by [10] which demonstrates that images of any resolution can be transformed to a fixed set of visual tokens while maintaining the semantic information, we propose to extract a V-POR by exploiting the intrinsic structure of the image-report pair and by the aid of the T-POR without extra human annotations. Specifically, we first introduce a visual pathology observation extractor (VPOE) which is a transformer-based architecture (adhering to a Self-Attention $\rightarrow$ Cross-Attention structure) with a sequence of learnable pathology query tokens $Q\in\mathbb{R}^{N_{q}\times D}$. $N_{q}$ denotes the number of query tokens. Query tokens, designed to extract specific V-PORs from localized visual tokens, first performs a self-attention mechanism to capture contextual relationships within themselves. Then, they interact with these localized visual tokens through a cross-attention mechanism within the VPOE to align/extract the VPOs. Here, the query tokens serve as queries, while the localized visual tokens act as keys and values. The proposed PCMA module utilizes T-POR and contrastive learning to guide the model in exploiting the inherent structure of the image-report pair. This supervision aids in training the pathology query tokens to extract the most pertinent V-POR. Given the local visual tokens $U^{LI}_{i}$ and query tokens $Q$, the process to obtain the V-POR for $l^{th}$ transformer layer $O^{PI}_{i,l}$ in VDE is summarized by Equation 5. The upper right part of Figure 2 also illustrates this process.
@@
where $\mathrm{-Attn}$ refers to a vanilla attention mechanism and $Q_{0}=Q$. We consider the output of the last transformer layer in $VDE$ as the final visual pathological observation representation denoted by $O^{PI}_{i}$. After that, we enforce the alignment between the visual and textual POR. Nevertheless, there are no ground truth matching annotations between the visual and textual POR since the T-POR for one anatomical region can occur at any position in the report. Similarly to [2, 3], we adopt a cross-attention mechanism to investigate the matching between visual and textual PORs. Specifically, for the T-POR $O^{PI}_{i,j}$, we generate its corresponding cross-modal attended representation $C^{PI}_{i,j}$ via:
@@
Then, similar to ICMA, we apply the symmetric InfoNCE loss to maximise mutual information between a V-POR $O^{PI}_{i}$ and its corresponding visual cross-modal attended representations (V-CARs) $C^{PI}_{i}$. Pathological-level report-to-image alignment loss $\mathcal{L}^{I\leftarrow R}_{PCMA}$ is formulated as:
@@
$w^{\prime}_{i}=\frac{w_{i}}{\sum_{k}{w_{k}}},\quad\quad w_{i}=\sum_{j\in sent_{i}}{a_{j}}.$ (9)
@@ 3.5 Cross-modal Correlation Exploration
where $\bar{e}^{i}$ and $\bar{e}^{j}$ are the mean of $e^{i}$ and $e^{j}$ respectively. By calculating the covariance among each pair of patches (variables), we obtain a covariance matrix $\bm{\Sigma}^{i}\in\mathbb{R}^{N_{p}\times N_{p}}$ describing the correlation among different image patches in the $i^{th}$ sample. After that, to enforce that the model capture further fine-grain details and promote the cross-modal representation learning, we design a novel proxy task which requires it to predict this covariance matrix conditioned on the global report representation. We simply employ one linear layer as the CCP head to predict the covariance matrix: $\bm{H}^{i}=W_{h}\cdot\bm{U}^{GR}_{i}$ where $\bm{H}^{i}$ refers to the predicted covariance matrix for $i^{th}$ sample. After that, the Mean Squared Error (MSE) loss is used to supervise the learning of the cross-modal correlation prediction:
@@
The proposed objective CCE necessitates a significantly more fine-grained details understanding compared to the traditional MIM since it demands not only that the model comprehends the content within each image region from reports but also discerns the potential correlations among various image regions. The CCE objective uses the global report representation to predict the covariance matrix. Gradients from the CCE head backpropagate through the textual encoder. The global textual representation interacts with localized textual tokens via self-attention, and these localized textual tokens further interact with V-PORs through cross-attention, as specified in the PCMA module. As a result, both the VPO Extractor and the visual encoder—which generates the localized visual tokens—are also influenced by the CCE objective.
@@
Table 1: Results of semantic segmentation and image classification in a zero-shot setting (for classification), 1%, 10% and 100% training samples. The evaluation metric for CheXpert is AUC. The best results are highlighted in bold.
@@ 3.6 Overall Objective
Our proposed model is jointly trained with the three modules, i.e. the ICMA, PCMA and CCE modules, enforcing the framework to learn a more fine-grained, generalizable and pathologically discriminative medical visual representation. The overall training objectives is formulated as:
@@ 4.1 Pre-training Setup
Dataset: We follow most previous studies [2, 3, 8] to pre-train our PLACE on the MIMIC-CXR dataset [36] and adopt the same data pre-processing procedure as [8, 9]. The images of the lateral view are removed since the downstream tasks only contain the frontal view images. Reports are formed by concatenating the Finding and Impression sections. We remove samples with an empty report or less than three words, resulting in a roughly total of 217,000 image-report pairs.
Implementation Details: Following [2, 3, 8], we employ ResNet-50 [29] and ClinicalBERT [37] as the backbone of our image and report encoder. Note that our proposed method is model-agnostic and can be applied to various backbones such as a vision transformer and convolutional networks. We also report the results of adopting the vision transformer as the image encoder in classification tasks. Images are first resized to $256\times 256$ while maintaining the original size ratio with zero-padding for the smaller dimension, and then randomly cropped to $224\times 224$ during training. We adopt the AdamW [38] as the optimizer with a learning rate of 4e44e-4 and weight decay of 5e25e-2. The batch size is set to 128 on three A100-40G GPU cards. We train our model for $50$ epochs with an early stop mechanism that terminates the training without seeing a decrease in validation loss for more than $10$ epochs. Consistent with [8, 39], we set the softmax temperature $\tau_{1}$ and $\tau_{2}$ to $0.07$ and $0.10$ respectively. The split patch size $PS$ in CCP module is set to $32$ (determined by a small grid search in $\{16,28,32,56\}$, resulting in a $N_{p}$ of $49$. We set the number of pathological query tokens in Equation 8 to $12$. The loss weights $\lambda$ and $\beta$ are set to $0.5$ determined by a small grid search in $\{0.1,0.25,0.5,1\}$.
@@ 4.2 Downstream Tasks
Medical Semantic Segmentation. The SIIM Pneumothorax [40] and RSNA Pneumonia [41] datasets are used to assess the capability of our model for this task. Consistent with most previous works, we adopt the U-Net architecture and employ our pre-trained image encoder as the the encoder backbone (weight frozen), while fine-tuning the decoder using 1%, 10% and 100% training samples. Dice score [42] is selected as the evaluation metric.
@@
Medical Image Classification. We verify the effectiveness of our pre-trained image encoder for medical image classification on COVIDx [44] and CheXpert [45] datasets. Following prior work [8, 9, 3], we adopt linear probing which freezes the pre-trained image encoder and only trains the classification head on COVIDx dataset on three scenarios, i.e., 1%, 10% and 100% training samples. Additionally, we also gauge the zero-shot generalization capability of our model on CheXpert and COVIDx datasets. Same as  [8, 9], we also report the results of taking the vanilla ViT [30] as the image encoder.
@@
Medical Report Generation. We further investigate PLACE’s ability to comprehend cross-modal information on the MIMIC-CXR dataset through the task of report generation. We adopt two fundamental and typical architecture as our baseline (1) ST [46]: a visual extractor with transformer-based encoder-decoder architecture and (2) C2GPT2 [47]: a vision encoder-language decoder architecture. Note that the baselines are standard transformer-based models without any advanced techniques or additional data sources tailored specifically to MRG. In our approach, we employ the pretrained image encoder from PLACE as the visual extractor in ST and as the vision encoder in Res2GPT2, maintaining fixed weights while only optimizing other model components. The official data split is adopted. We evaluate the performance of the model by the commonly utilized metrics including BLEU [48], METEOR [49], ROUGE-L [50] and CIDEr [51].
@@
Prec @@ 1 Prec @@ 2 Prec @@ 5 Prec @@ 10